# MapCoder-Lite: Squeezing Multi-Agent Coding into a Single Small LLM

### v1

@@ 3.2 Supervisor-Guided Planning and Coding
Observations. Even after the role-specific fine-tuning on strong LLM trajectories, the 7B planning agent still emits fragmentary execution flows, omitting edge cases, misordering steps, or drifting from the prescribed outline. Also, the coding agent continues to generate programs that compile yet collapse under the full unit-test suite. These local errors propagate: an incomplete plan starves the coder of crucial hints, and a superficially correct program deceives the debugger until final evaluation. Table 2 confirms that this trajectory-only adaptation closes part of the gap but suffers from a noticeable accuracy loss. In contrast, Qwen2.5-32B-Instruct and DeepSeek-V3 produce structurally sound plans and largely correct first-pass implementations, indicating that stronger models possess global cues that the smaller model has not yet internalised.
Challenges. Despite the availability of agent-wise strong-model trajectories, two obstacles remain for high-quality fine-tuning. First, a locally tuned agent lacks global awareness: improving its own output does not guarantee end-to-end success unless that output is judged in context. Second, the capacity gap hinders direct imitation: a 7B model exposed to large-scale models’ plans and coding results may underfit or copy surface patterns without learning the deeper reasoning (phenomenon called capacity gap Bansal et al. (2024); Xu et al. (2025)).
Methods. We propose a supervisor-guided refinement pipeline that supplies global feedback without enlarging the runtime model. As shown in Fig. 4, a 7B MapCoder first produces a complete retrieval→planning→coding trajectory. If the resulting program fails its unit tests, the entire trajectory is forwarded to a high-capacity supervisor LLM (DeepSeek-V3). The supervisor analyses the trace, pinpoints the agent primarily responsible for the failure, and issues concise, role-specific feedback, and only that agent regenerates its output. The revised trajectory is re-tested; once all tests pass, the final plan–code pair is added to the fine-tuning corpus.
Supervisor-guided refinement is applied selectively to trajectories where the strong LLM succeeds but the 7B model fails, minimizing generation overhead. The supervisor operates only during data generation, keeping inference lightweight. We store just the corrected agent input–output pairs, omitting the feedback itself, so the 7B model learns solely from information available at runtime. Crucially, every example in this corpus is produced and execution-validated by the 7B model, eliminating concerns about a capacity gap. Iterating over thousands of problems yields a dataset that aligns planners and coders with end-to-end success. Table 2 confirms that this strategy delivers a substantial accuracy boost.
FT Method
Trainable
Parameters
Required Memory
for Training (GB)
Accuracy
@@
3.3 Multi-Agent LoRA Fine-Tuning
To enable role-specific specialization in our multi-agent framework, we adopt an agent-wise LoRA approach. All four agents—retrieval, planning, coding, and debugging—share a frozen 7B base model, each paired with its own lightweight LoRA. These adapters are fine-tuned independently using pass-based-filtered and supervisor-guided trajectory distillation. This setup significantly reduces GPU memory usage during fine-tuning and supports modular training and deployment, allowing each agent to be updated without duplicating the full model. As shown in Table 3, LoRA significantly reduces the number of trainable parameters and memory footprint, while achieving even higher accuracy than full fine-tuning. We attribute this accuracy gain in part to LoRA’s implicit regularization, which steers the model toward role-specific behaviors without overwriting core pretrained knowledge—a finding consistent with prior work Hu et al. (2021); Houlsby et al. (2019).
4 Experimental Settings
@@
Models. All agent-wise fine-tuning experiments use Qwen2.5-7B-Instruct as the base model. For comparison, we include larger models—Qwen2.5-14B-Instruct and Qwen2.5-32B-Instruct, serving as upper bounds to assess the effectiveness of our lightweight strategy.
Benchmarks. We evaluate models on three competitive programming benchmarks: xCodeEval Khan et al. (2024), APPS Hendrycks et al. (2021), and CodeContests Li et al. (2022). For xCodeEval, we use the 106 problems from the compact program synthesis split. For CodeContests, we evaluate all 165 official test problems. We randomly select 150 test-set problems for APPS, following MapCoder Islam et al. (2024). Each benchmark includes natural language problem descriptions with sample I/O and is evaluated using hidden unit tests. We also optionally report results on HumanEval and MBPP to assess generalization to function-level tasks.
Training Configuration. Each agent is fine-tuned using LoRA adapters with rank 32 applied to the linear modules for query, key, value, and output projection. Training is implemented with Hugging Face Transformers and PEFT libraries HuggingFace , using a learning rate of 2e-5, gradient accumulation of 16, and 3 training epochs. Hyperparameters were tuned individually per agent to ensure stable and consistent performance. All training experiments were conducted on a single NVIDIA A100 80GB GPU.
Evaluation Setting. We report pass@1 accuracy as the primary metric, reflecting whether the model’s top-1 output passes all test cases. All outputs are generated using greedy decoding for reproducibility and consistency with baseline evaluation.
5 Experimental Results
@@ 5.1 Effectiveness of Agent-wise Fine-Tuning
Table 4 compares the performance of a small LM (Qwen2.5-7B) across various single-agent prompt optimization methods—direct prompting, chain-of-thought (CoT) Wei et al. (2023), self-planning Jiang et al. (2024b), and analogical prompting Yasunaga et al. (2024)—as well as multi-agent approaches: MapCoder and our proposed MapCoder-Lite. Without fine-tuning, MapCoder shows only marginal gains over single-agent methods and even underperforms Direct prompting on xCodeEval (13.21% vs. 18.87%), despite using four specialized agents. This highlights the limitations of small LMs in multi-agent setups, where poor coordination and frequent format violations (as indicated by numerous format failures) prevent reliable end-to-end performance.
In contrast, MapCoder-Lite, with agent-wise fine-tuning, achieves substantial improvements across all benchmarks. On xCodeEval, accuracy more than doubles to 28.30%, outperforming all prompting baselines. Similar gains are seen on APPS (6.00% → 8.00%) and CodeContests (6.06% → 13.33%). Notably, fine-tuning eliminates all format failures, indicating that even small LMs can produce structurally consistent outputs when role-specialized. These results emphasize the critical role of agent-wise fine-tuning in realizing the full potential of small LMs in multi-agent frameworks.
@@ 5.2 Comparison with Larger Models
5.3 Ablation Study
Contribution of Individual Agent Fine-Tuning. Table 6 shows an ablation study evaluating the impact of fine-tuning each agent in MapCoder on xCodeEval. Fine-tuning the retrieval agent alone improves accuracy from 13.21% to 18.87% and reduces format failures from 29 to 12, indicating better structural consistency. Adding the planning agent further boosts accuracy to 26.42% and eliminates format errors entirely, increasing both non-debug passes (17 → 18) and debug-assisted completions (3 → 10). Fine-tuning the coding agent (without debugging) raises non-debug passes to 24 but reduces debug-assisted passes to 2, since MapCoder only invokes the debugger when initial code fails sample I/O tests. Stronger code may bypass debugging entirely, leaving some errors uncaught. Conversely, fine-tuning the debugging agent (without coding) increases accuracy to 27.36%, with 12 debug-assisted completions—showing that an effective debugger can recover from weaker initial code. When all four agents are fine-tuned, overall accuracy peaks at 28.30%, with a strong balance between direct (24) and debug-assisted (6) completions. These results demonstrate that each agent addresses distinct failure modes and that coordinated fine-tuning is essential for maximizing end-to-end performance.
@@ 5.3 Ablation Study
Multi-Agent Tuning with Coder Model. We evaluate whether multi-agent code generation can be enhanced using a code-specialized language model (a.k.a. Coder model), such as Qwen2.5-Coder-7B-Instruct Hui et al. (2024). As shown in Table 8, applying the multi-agent framework (MapCoder) and our fine-tuning strategy (MapCoder-Lite) to the Coder model does improve its performance, demonstrating that even domain-specialized models benefit from structured, role-specific tuning. However, despite these gains, the Coder-based pipeline still lags significantly behind our baseline model, Qwen2.5-7B-Instruct, both before and after fine-tuning. We attribute this gap to the Coder model’s limited general reasoning ability, which is critical for handling the diverse roles in multi-agent workflows, including retrieval, planning, and debugging, in addition to coding. In contrast, the general-purpose Qwen2.5-7B-Instruct model more effectively supports multi-agent coordination, and with our memory-efficient fine-tuning (MapCoder-Lite), achieves even higher end-to-end accuracy.
Model Method Accuracy (%)
Qwen2.5-Coder-7B-Instruct MapCoder 5.66
MapCoder-Lite 15.09
Qwen2.5-7B-Instruct MapCoder 11.32
MapCoder-Lite 22.64
Table 8: Comparison of MapCoder and MapCoder Lite using Qwen2.5-Coder-7B-Instruct. Accuracy is measured only up to the coding stage, without debugging.
5.4 Case Study
To qualitatively assess the impact of agent-wise fine-tuning, we examine representative before-and-after examples for each agent. As illustrated in Appendix B, fine-tuning leads to notable improvements in algorithm retrieval, plan coverage, code correctness, and bug resolution. These examples highlight how each agent benefits from role-specific supervision, supporting the effectiveness of our fine-tuning strategy.
@@ 6 Conclusion
We introduced MapCoder-Lite, a multi-agent coding framework built on a 7B model with lightweight LoRA adapters. Through trajectory distillation, supervisor-guided refinement, and agent-wise specialization, it achieves competitive performance with lower cost. Our results show that small models, when carefully fine-tuned, can match the reliability and accuracy of larger systems.
@@
Our framework relies on distilled trajectories from strong models like DeepSeek-V3 and Qwen2.5-32B, yet the fine-tuned 7B model does not fully replicate their performance. Rather than indicating a ceiling, this gap highlights opportunities to further enhance small models—potentially through architectural extensions, better training objectives, or reinforcement-based learning from distillation feedback.
In addition, the multi-agent structure introduces a wide design space for fine-tuning, as each agent has a distinct role and optimization goal. While we adopted a uniform tuning recipe across agents, this modularity also enables targeted specialization and co-training strategies in future work, suggesting that more adaptive tuning could unlock even greater performance.
@@ Appendix A Details of Supervisor
Fig. 5 and Fig. 6 illustrate the operation of the supervisor used in our data curation pipeline. The supervisor receives the full execution trajectory of a problem, including the problem description, intermediate agent outputs (e.g., algorithm explanation, plan, code), and the result of test case execution.
@@
As shown in Fig. 5, the input prompt explicitly instructs the supervisor to analyze the trajectory, determine which agent is responsible for the failure, and provide natural language feedback targeting that agent. The goal is to isolate the error at the correct stage (retrieval, planning, or coding), and produce actionable guidance that can be used to re-generate only the faulty component.
@@
Fig. 6 shows a typical response from the supervisor. It correctly attributes the failure to the planning agent, explains why the plan is insufficient, and suggests how it should be modified. This output is then used to re-invoke the corresponding agent with additional guidance, creating high-quality training data without requiring human annotation.
Appendix B Improvements After Fine-Tuning

### v2

@@ 5.2 Supervisor-Guided Planning and Coding
Even when fine-tuned on strong-model trajectories, 7B agents struggle to achieve robust multi-stage reasoning. One challenge is the lack of global awareness, where agents trained in isolation fail to account for downstream dependencies, an issue previously discussed in Section 4. Another is the capacity gap Bansal et al. (2024); Xu et al. (2025): while strong LLMs produce coherent and mostly correct outputs, small models tend to mimic surface forms without learning the underlying reasoning. These limitations underscore the need for supervision that supports both global coordination and deeper abstraction.
To address these issues, we propose a supervisor-guided refinement pipeline that supplies global feedback without enlarging the runtime model. As shown in Fig. 5, a 7B MapCoder first produces a complete retrieval→planning→coding trajectory. If the resulting program fails its unit tests, the entire trajectory is forwarded to a high-capacity supervisor LLM (DeepSeek-V3). The supervisor analyzes the trace, identifies the agent primarily responsible for the failure, and issues concise, role-specific feedback. Only the selected agent then regenerates its output. The revised trajectory is re-tested; once all tests pass, the final plan–code pair is added to the fine-tuning corpus.
Supervisor-guided refinement is applied selectively to trajectories where the strong LLM succeeds but the 7B model fails, minimizing generation overhead. The supervisor operates only during data generation, keeping inference lightweight. We store only the corrected agent input–output pairs, omitting the feedback itself, so the 7B model learns solely from information available at runtime. Crucially, every example in this corpus is produced and execution-validated by the 7B model, eliminating concerns about a capacity gap. Iterating over thousands of problems yields a dataset that better aligns planners and coders with end-to-end success. Table 1 confirms that this strategy delivers a substantial accuracy boost.
5.3 Multi-Agent LoRA Fine-Tuning
FT Method Trainable Parameters Required Memory for Training (GB) Accuracy (%)
@@ 5.3 Multi-Agent LoRA Fine-Tuning
Figure 6: Training and validation loss for the Planning agent using full fine-tuning (FFT) and LoRA. LoRA shows higher training loss but maintains lower and more stable validation loss, suggesting better generalization.
To enable role-specific specialization, we adopt an agent-wise LoRA strategy: all four agents share a frozen 7B backbone, each with independent rank-32 LoRA adapters fine-tuned on role-specific data. Unlike prior approaches that fine-tune the entire model across tasks Shen et al. (2024); Zeng et al. (2023), our method isolates agent behavior at the adapter level while preserving a unified backbone.
@@
We empirically show that LoRA serves effectively as a modular control layer in multi-agent code generation, delivering both accuracy and efficiency gains (Table 2). We attribute this to LoRA’s implicit regularization Hu et al. (2021); Houlsby et al. (2019), which promotes task-specific behavior without overwriting pretrained knowledge. In multi-agent settings with complex, structured I/O, full fine-tuning often overfits to superficial patterns, weakening general reasoning. As shown in Fig. 6, full fine-tuning achieves lower training loss but increasing validation loss, whereas LoRA maintains lower and more stable validation loss, indicating better generalization and adaptability across roles.
6 Experimental Settings
Models. All agent-wise fine-tuning experiments use Qwen2.5-7B-Instruct as the base model. For comparison, we include larger variants—Qwen2.5-14B-Instruct and Qwen2.5-32B-Instruct—which serve as upper bounds for evaluating the effectiveness of our lightweight strategy. To assess whether a code-pretrained backbone improves performance, we additionally test Qwen2.5-Coder-7B-Instruct. We also apply our method to Llama3.1-8B and CodeLlama-7B to demonstrate generalizability beyond the Qwen family.
Benchmarks. We evaluate models on three competitive programming benchmarks: xCodeEval Khan et al. (2024), APPS Hendrycks et al. (2021), and CodeContests Li et al. (2022). For xCodeEval, we use the 106 problems from the compact program synthesis split. For CodeContests, we evaluate all 165 official test problems. For APPS, we randomly select 150 test-set problems, following MapCoder Islam et al. (2024). Each benchmark provides natural language problem descriptions with sample I/O and is evaluated using hidden unit tests. We additionally report results on HumanEval and MBPP to assess generalization to function-level tasks.
Fine-Tuning Setup. Each agent is fine-tuned using LoRA adapters with rank-32, applied to the linear layers for query, key, value, and output projections. We use a learning rate of 2e-5, gradient accumulation of 16, and train for three epochs, with hyperparameters tuned per agent to ensure stability. All training runs are conducted on a single NVIDIA A100 80GB GPU. The cost of data curation prior to fine-tuning is summarized in Table 8 (Section 7.5).
Evaluation Setting. We report pass@1 accuracy using greedy decoding to ensure reproducibility and align with baseline evaluations.
7 Results
@@ 7.1 Effectiveness of Agent-wise Fine-Tuning
Table 3 compares the performance of Qwen2.5-7B under various prompting strategies—direct prompting, CoTWei et al. (2023), self-planningJiang et al. (2024b), analogical Yasunaga et al. (2024)—and multi-agent setups. Without fine-tuning, MapCoder yields only marginal improvements and even underperforms direct prompting on xCodeEval (13.21% vs. 18.87%), underscoring the difficulty of coordinating small models in multi-agent pipelines due to format errors and poor inter-agent coordination.
In contrast, MapCoder-Lite with agent-wise fine-tuning achieves substantial improvements across all benchmarks. On xCodeEval, accuracy more than doubles to 28.30%, outperforming all prompting baselines. Similar gains are observed on APPS (6.00% → 8.00%) and CodeContests (6.06% → 13.33%). Notably, fine-tuning eliminates all format failures, indicating that even small LMs can produce structurally consistent outputs when role-specialized. These results highlight the critical role of agent-wise fine-tuning in realizing the full potential of small LMs in multi-agent frameworks, with qualitative examples in Appendix E further illustrating improvements across all roles.
@@ 7.2 Comparison with Larger Models
7.3 Ablation Study
Contribution of Individual Agent Fine-Tuning. Table 5 presents an ablation study of agent-wise fine-tuning on xCodeEval. Tuning the retrieval agent alone raises accuracy from 13.21% to 18.87% and cuts format failures from 29 to 12. Adding the planning agent further improves accuracy to 26.42% and eliminates all format errors, boosting both non-debug and debug-assisted completions. Fine-tuning the coding agent (without debugging) increases non-debug passes to 24 but reduces debug-assisted ones, as stronger initial code often bypasses the debugger. In contrast, tuning the debugging agent (without coding) improves recovery from weak code, achieving 27.36% accuracy and 12 debug-assisted completions. Full fine-tuning of all agents yields the best performance (28.30%), confirming that each agent contributes uniquely and coordinated tuning is critical for optimal results.
@@ 7.3 Ablation Study
Retrieval Planning Coding xCodeEval (%)
32 32 32 22.64
8 32 32 18.87
32 8 32 20.75
32 32 8 21.70
8 8 8 21.70
64 32 32 16.98
32 64 32 20.75
32 32 64 20.75
64 64 64 16.98
Table 6: Effect of LoRA rank on multi-agent performance. The three leftmost columns indicate the LoRA ranks applied to the retrieval, planning, and coding agents, respectively.
@@
Impact of LoRA Rank. We study the effect of LoRA rank on multi-agent performance to justify the configuration used in this work. The rank-32 setting was selected based on empirical evaluation rather than by convention. As shown in Table 6, rank-32 achieves the highest xCodeEval accuracy (22.64%) among all tested configurations. Lower-rank settings (e.g., $(8,8,8)$) underperform despite being parameter-efficient, while higher-rank configurations (e.g., 64) do not yield further gains and can degrade performance. These results suggest that multi-agent performance depends on interactions between agent roles rather than parameter count alone. Overall, rank-32 offers the best balance between accuracy and efficiency, introducing only about 3% additional trainable parameters relative to the backbone.
Backbone Selection. Choosing the right backbone is critical for multi-agent code generation, where each agent performs a distinct, reasoning-intensive role. Among SLM candidates, we compare a general-purpose model (Qwen2.5-7B-Instruct Qwen et al. (2025)) and a code-specialized variant (Qwen2.5-Coder-7B-Instruct Hui et al. (2024)) for use in MapCoder. Despite similar sizes, the general-purpose model achieves higher accuracy and format adherence. Agent-wise ablations further show that replacing even one agent with its coder counterpart degrades performance, suggesting that the coder model lacks the contextual reasoning and adaptability needed for role-specific tasks. These results support our choice of the general-purpose model as the unified backbone for MapCoder-Lite (see Appendix B).
7.4 Generalization and Robustness
@@
Model Compatibility. Our method is architecture-agnostic and applicable to LoRA-supported backbones. Beyond the Qwen2.5 family, MapCoder-Lite improves accuracy and reduces format failures on Llama3.1-8B and CodeLlama-7B across xCodeEval and HumanEval (Appendix D). We also evaluate MapCoder-Lite on the recent reasoning-oriented model Qwen3-4B; despite strong single-agent prompting performance, stable multi-agent behavior is only achieved after agent-wise fine-tuning, with detailed results reported in Appendix C.
7.5 Data Curation Cost and Statistics
Table 8 summarizes the dataset size, resource usage, and data generation time for each agent. In particular, we curate approximately 2.3k execution-verified trajectories for the retrieval agent, 1.1k each for the planning and coding agents, and 4.2k for the debugging agent. We used 2×A100 GPUs with vLLM serving to run Qwen2.5-32B and Qwen2.5-7B models, and partially relied on the DeepSeek API as supervisor and debugging agent, incurring \$20–30. Data generation is performed only once, and the resulting datasets are reusable across models and experiments. The strong supervisor model is invoked only in failure cases to provide brief corrective feedback, rather than full trajectory generation, keeping strong-model involvement infrequent and tightly bounded to training-time supervision. Retrieval and debugging took longer due to longer outputs; debugging in particular depends on planning and coding outputs, requiring upstream completion. While data was generated per agent during development, the pipeline can be streamlined into a single pass for faster future use.
Retrieval Planning, Coding Debugging
Dataset size 2.3k 1.1k (each) 4.2k
Time 2–3 days 16 hours 4–5 days
GPUs 2 $\times$ A100 2 $\times$ A100 2 $\times$ A100
Table 8: Dataset size, data generation time, and computational resource usage for data curation for each agent.
@@ 8 Conclusion
We introduced MapCoder-Lite, a multi-agent coding framework built on a 7B model with lightweight LoRA adapters. Through trajectory distillation, supervisor-guided refinement, and agent-wise specialization, it achieves competitive performance at substantially lower cost. Our results demonstrate that small language models, when carefully fine-tuned, can match the reliability and accuracy of much larger systems in multi-agent code generation.
Acknowledgments
This work was supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. RS-2025-00561961). This work was also supported by the Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (No. RS-2025-02214497, Development of Low-Level Optimization Program API Technology for AI Semiconductors).
@@ Acknowledgments
Our framework relies on distilled trajectories from strong models such as DeepSeek-V3 and Qwen2.5-32B, yet the fine-tuned 7B model does not fully replicate their performance. Rather than indicating a hard performance ceiling, this gap highlights opportunities to further enhance small models, potentially through architectural extensions or improved training objectives. In addition, the multi-agent structure introduces a broad design space for fine-tuning, as each agent has a distinct role and optimization objective. While we adopt a uniform tuning recipe across agents in this work, more adaptive tuning could further improve performance.
@@ Appendix A Details of Supervisor
Fig. 7 and Fig. 8 illustrate the operation of the supervisor in our data curation pipeline. The supervisor receives the full execution trajectory of a problem, including the problem description, intermediate agent outputs (e.g., algorithm explanation, step-by-step plan, and code), and the result of test case execution.
@@
As shown in Fig. 7, the input prompt instructs the supervisor to analyze the trajectory, determine which agent is responsible for the failure, and provide natural language feedback targeting that agent. The goal is to isolate the error at the correct stage (retrieval, planning, or coding), and produce actionable guidance that enables re-generation only the faulty component.
@@
Fig. 8 shows a representative response from the supervisor. It correctly attributes the failure to the planning agent, explains why the plan is insufficient, and suggests how it should be modified. This output is then used to re-invoke the corresponding agent with additional guidance, creating high-quality training data without human annotation.
Appendix B Backbone Model Comparison: General-Purpose vs. Coder
To investigate the impact of backbone model selection in multi-agent code generation, we conducted a series of controlled experiments comparing general-purpose and code-specialized models. Specifically, we tested Qwen2.5-7B-Instruct (general-purpose) and Qwen2.5-Coder-7B-Instruct (code-specialized) under various configurations of the MapCoder and MapCoder-Lite pipelines. The results show that general-purpose models are more robust across all agent roles and better suited for both zero-shot and fine-tuned multi-agent setups.
Benchmark Performance (Zero-shot). Table 9 shows the performance of MapCoder using each backbone model. The general-purpose variant consistently achieves higher accuracy and fewer format failures, highlighting its superiority in multi-agent coordination and instruction-following tasks.
Benchmark Metric 7B 7BCoder
[…]
Table 9: Comparison of Qwen2.5-7B-Instruct and Qwen2.5-Coder-7B-Instruct across benchmarks.
Agent-wise Ablation (Zero-shot). As shown in Table 10, replacing any agent in the general-purpose pipeline with its coder counterpart leads to a consistent accuracy drop, most notably in the planning and coding agents. Even for the coding role, where the coder model is expected to excel, performance decreases—highlighting the importance of upstream integration and general reasoning.
Retrieval Planning Coding Debugging xCodeEval (%)
7B 7B 7B 7B 13.21
7BCoder 7B 7B 7B 12.26
7B 7BCoder 7B 7B 10.38
7B 7B 7BCoder 7B 10.38
7B 7B 7B 7BCoder 12.26
Table 10: Ablation of 7B vs. 7B Coder in retrieval, planning, coding, and debugging agents on the xCodeEval benchmark.
Training Loss after Fine-tuning. We fine-tuned both models using identical LoRA settings across all agents. As summarized in Table 11, the coder model converges to higher final loss values, suggesting a poorer fit to role-specialized training data. This pattern is consistent across all agents.
Model Retrieval Planning Coding Debugging
Qwen2.5-7B-Instruct 0.34 0.25 0.08 0.33
Qwen2.5-Coder-7B-Instruct 0.36 0.35 0.13 0.34
Table 11: Final training loss after LoRA fine-tuning for each agent using general-purpose vs. coder-specific models.
Ablation after LoRA Fine-tuning. Table 12 presents accuracy when substituting each LoRA-finetuned agent with its coder-based variant. Accuracy declines in all settings, and the fully coder-based pipeline achieves only 18.87%, compared to 28.30% for the general-purpose variant. These results indicate that even after fine-tuning, general-purpose models better support the multi-agent workflow.
Retrieval Planning Coding Debugging xCodeEval (%)
7BFT 7BFT 7BFT 7BFT 28.30
7BCoderFT 7BFT 7BFT 7BFT 21.70
7BFT 7BCoderFT 7BFT 7BFT 23.58
7BFT 7BFT 7BCoderFT 7BFT 24.53
7BFT 7BFT 7BFT 7BCoderFT 23.58
7BCoderFT 7BCoderFT 7BCoderFT 7BCoderFT 18.87
Table 12: xCodeEval accuracy when substituting each agent with a LoRA-finetuned coder-specific model (7BCoderFT), compared against the general-purpose baseline (7BFT).
Appendix C Evaluation on Recent Reasoning Models (Qwen3)
We evaluate MapCoder-Lite on Qwen3-4B, a recently released small language model that significantly improves single-agent reasoning and outperforms Qwen2.5-7B under direct prompting. As shown in Table 13, Qwen3-4B achieves strong direct-prompting performance (22.64% without thinking and 26.42% with thinking). However, this improvement does not directly translate into reliable multi-agent behavior. When deployed in the MapCoder pipeline without fine-tuning, Qwen3-4B attains only 3.77% accuracy, exhibiting the same instability observed with earlier backbones.
Applying MapCoder-Lite mitigates this issue. After fine-tuning the retrieval, planning, and coding agents, Qwen3-4B achieves 30.19% accuracy, surpassing the fine-tuned Qwen2.5-7B despite using fewer parameters. These results indicate that recent advances in single-agent reasoning alone are insufficient for stable multi-agent workflows, and that targeted agent-wise specialization remains essential. Moreover, the larger gains observed with Qwen3-4B suggest that MapCoder-Lite scales positively with backbone capability, effectively activating role-specific behaviors that are not induced by prompting alone.
Method Qwen2.5-7B-Instruct Qwen3-4B
Direct (non-thinking) 18.87 22.64
Direct (thinking) 26.42
MapCoder (RPC) 11.32 3.77
MapCoder-Lite (RPC) 22.64 30.19
Table 13: xCodeEval performance of Qwen2.5-7B and Qwen3-4B under direct prompting and MapCoder pipelines. RPC denotes the retrieval–planning–coding agents.
Appendix D Generalization Across Model Families
While our main experiments used the Qwen series due to their strong coding performance on HumanEval and MBPP during development, our method remains architecture-agnostic. Since MapCoder-Lite applies LoRA through the PEFT library, it can be seamlessly used with any compatible model.
To test this generalizability, we evaluated both MapCoder and MapCoder-Lite using Llama3.1-8B-Instruct and CodeLlama-7B-Instruct-hf as backbones. As shown in Table 14, both models performed poorly without fine-tuning on competitive programming tasks. For example, Llama3.1-8B scored 0.00% with 94 format failures on xCodeEval. After applying MapCoder-Lite, Llama3.1-8B achieved 16.04% accuracy with zero format failures, demonstrating the adaptability of our pipeline. In contrast, CodeLlama-7B showed limited improvement on xCodeEval but responded well on function-level benchmarks such as HumanEval.
Model Benchmark MapCoder MapCoder-Lite
Accuracy (%) Accuracy (%)
Llama3.1-8B-Instruct xCodeEval 0.00 16.04
CodeLlama-7B-Instruct xCodeEval 0.94 2.83
CodeLlama-7B-Instruct HumanEval 10.98 45.12
Table 14: Comparison of MapCoder and MapCoder-Lite on Llama3.1-8B and CodeLlama-7B models on xCodeEval and HumanEval benchmarks.
Additionally, we compare MapCoder-Lite with various prompting strategies using the Llama3.1-8B backbone. As summarized in Table 15, MapCoder underperforms direct prompting baselines, whereas MapCoder-Lite substantially improves accuracy and surpasses all prompting methods. These results reinforce that multi-agent pipelines require targeted fine-tuning to be effective: without it, they underperform simpler approaches, while with it, they become competitive and scalable.
Method Accuracy (%)
Direct 6.60
CoT 4.72
Self-planning 4.72
Analogical 5.66
Multi-agent (MapCoder) 0.00
Multi-agent + FT (MapCoder-Lite) 16.04
Table 15: Accuracy comparison between MapCoder-Lite and other prompting baselines on xCodeEval using Llama3.1-8B-Instruct.
Appendix E Improvements After Fine-Tuning